Suppose you toss a half-dollar coin n times. How large must n be to guarantee that your probability of getting heads at least once is better than 0.99?
Answer #1

P(at least one head) =1-P(no head in n toss) =1-(1/2)n >=0.99

(1/2)n <=0.01

taking log on both sides

n >=ln(0.01)/ln(0.05)

n>=6.64

n.=7